# Fine-Tuning Vision-Language-Action Models: Optimizing Speed and Success

### v1

@@ II Related Work
Prior works have leveraged language and vision foundation models to enhance robotic capabilities, utilizing them as pretrained visual representations that accelerate robotic policy learning [29, 32, 30, 19, 31], for object localization in robotics tasks [9, 45], and for high-level planning and reasoning [1, 17, 42, 16, 43, 18, 6]. More recently, researchers have explored fine-tuning vision-language models (VLMs) to directly predict low-level robotic control actions, producing “vision-language-action” models (VLAs) [4, 33, 23, 22, 7, 15, 8, 50, 55, 51, 3, 2], which have demonstrated effective generalization to out-of-distribution test conditions and unseen semantic concepts. These works focus primarily on model development, while we focus on developing a recipe for fine-tuning such models, justifying individual design decisions with insights that we gain from our empirical analysis.
@@ IV-B Implementing Alternative Design Components
A crucial implementation detail is the choice of what represents a “feature” for modulation in vision transformers. While one might naturally consider treating individual patch embeddings as features to be modulated, we find that this approach results in poor language following. Instead, drawing from how FiLM operates in convolutional networks, where modulation applies spatially-agnostically by scaling and shifting entire feature maps, we apply each element of $\mathbf{\gamma}$ and $\mathbf{\beta}$ to the corresponding hidden unit across all visual patch embeddings so that $\mathbf{\gamma}$ and $\mathbf{\beta}$ influence all patch embeddings. Concretely, this makes $\mathbf{\gamma}$ and $\mathbf{\beta}$ $D_{ViT}$-dimensional vectors, where $D_{ViT}$ is the number of hidden dimensions in visual patch embeddings.
@@ V-A LIBERO Experimental Setup
TABLE I: LIBERO task performance results. Success rates (SR) across LIBERO benchmark task suites [25]. Results include policies fine-tuned from pretrained base models (Octo, DiT Policy, Seer, $\pi_{0}$), models trained from scratch (Diffusion Policy, Seer (scratch), MDT), and OpenVLA variants using different fine-tuning design decisions: parallel decoding (PD), action chunking (AC, chunk size $K=8$ timesteps), and continuous actions with L1 regression (Cont-L1) or with diffusion (Cont-Diffusion). OpenVLA results from this work are averaged over 500 trials for each task suite (10 tasks × 50 episodes). Our full OpenVLA-OFT method with additional inputs achieves state-of-the-art 97.1% average success rate. Baseline results are from the original papers, except for Diffusion Policy, Octo, and the original OpenVLA policies, whose results are reported by Kim et al. [22]. Bold and underlined values indicate best and second-best performance.
@@
OpenVLA (fine-tuned) [22] 84.7 88.4 79.2 53.7 76.5
OpenVLA (fine-tuned) + PD&AC 91.3 92.7 90.5 86.5 90.2
OpenVLA (fine-tuned) + PD&AC, Cont-L1 96.2 98.3 96.2 90.7 95.3
@@ V-B LIBERO Task Performance Comparisons
TABLE II: LIBERO inference efficiency results. Action generation throughput and latency for 7-dimensional actions, averaged over 100 queries on an NVIDIA A100 GPU. Each query processes a 224 × 224 px image and a LIBERO task command (“pick up the alphabet soup and place it in the basket”). We compare OpenVLA variants with parallel decoding (PD), action chunking ($K=8$ timesteps), and continuous actions using L1 regression or diffusion objectives. The last row adds wrist camera image and robot state inputs. Bold and underlined values show best and second-best performance.
Throughput (Hz) $\uparrow$ Latency (Sec) $\downarrow$
OpenVLA 4.2 0.240
+ PD 15.9 0.063
+ PD&AC 108.8 0.074
+ PD&AC, Cont-L1 109.7 0.073
+ PD&AC, Cont-Diffusion 10.1 0.792
@@
       + Additional Inputs (wrist img, proprio) 71.4 0.112
@@ V-C LIBERO Inference Efficiency Comparisons
Results in Table II show that parallel decoding reduces latency and increases throughput by 4$\times$ by replacing 7 sequential forward passes through the decoder portion of the policy with a single pass. Adding action chunking ($K=8$) increases latency by 17% due to longer attention sequences in the decoder, but when combined with parallel decoding, it dramatically improves throughput, achieving a 26$\times$ speedup over baseline OpenVLA. The continuous actions variant with L1 regression shows negligible difference in efficiency since the additional MLP action head adds minimal computational cost compared to the base model. The diffusion variant requires 50 denoising steps and thus suffers from 3$\times$ greater latency. However, it still achieves 2$\times$ higher throughput than baseline OpenVLA through parallel decoding and chunking. This means that despite longer pauses between action chunks, the diffusion variant still completes robot episodes more quickly than the original autoregressive variant.
@@ V-D Model Input-Output Flexibility
Evaluating these policies with additional inputs on the LIBERO benchmark reveals further improvements in average success rate across all task suites (Table I). Notably, our enhanced fine-tuned OpenVLA policies outperform even the best fine-tuned $\pi_{0}$ policies [3, 36]—which benefit from a base model with larger-scale pretraining and a more sophisticated learning objective (flow matching [24])—as well as the Multimodal Diffusion Transformer (MDT) [38]. Even with a simpler base model pretrained on less data than more recent VLAs, we find that our alternative VLA adaptation design decisions empower fine-tuned OpenVLA policies to establish a new state of the art on the LIBERO benchmark.
@@ V-F Additional Experiments
Given that the alternative fine-tuning formulation, along with additional model inputs and outputs, induces a distribution shift between the base VLA’s pretraining and fine-tuning, one reasonable question is whether the base VLA’s pretrained representations are helpful and have any influence on the results we have reported. We conduct an ablation study in Appendix A-G to address this question, ablating the OpenVLA pretraining phase and directly fine-tuning the underlying pretrained VLM with the OFT recipe. As shown in Table XIV, the base OpenVLA pretrained representations are indeed still beneficial for robotic policy learning, as removing them leads to a 5.2% drop in average success rate (absolute) in our LIBERO evaluation suite.
@@ VI-B Methods in Comparison
Figure 4: ALOHA task performance results. Comparison between policies trained from scratch (ACT, Diffusion Policy) and fine-tuned VLAs (RDT-1B, $\pi_{0}$, OpenVLA-OFT+) across four ALOHA manipulation tasks. OpenVLA-OFT+ enhances the base model with parallel decoding, action chunking, continuous actions, L1 regression, and FiLM [35] for language grounding. Fine-tuned VLAs consistently outperform from-scratch methods, with OpenVLA-OFT+ achieving the highest average performance.
@@ Acknowledgments
We thank the Stanford Center for Research on Foundation Models (CRFM) and Stanford Institute for Human-Centered AI (HAI) for providing computational resources that supported this research. Toyota Research Institute (TRI) provided funds to assist the authors with their research, but this article solely reflects the opinions and conclusions of its authors and not TRI or any other Toyota entity. This work was also in part supported by the Robotics and AI Institute and ONR grant N00014-22-1-2621. We thank Physical Intelligence for providing beta access to the $\pi_{0}$ model used in our ALOHA robot evaluations. We also thank Dan Fu, Karl Pertsch, and Alec Lessing for insightful discussions which contributed to the development of this work.
@@ A-D OpenVLA-OFT Hyperparameters and Training Details
OpenVLA-OFT+ training details for ALOHA. Hyperparameters for OpenVLA-OFT+ training on ALOHA tasks (with FiLM in the augmented OFT+ recipe) are shown in Table V. We maintain the same convergence criterion as in the LIBERO experiments (training until mean noramlized L1 loss falls below 0.01) and similar learning rate decay strategy (again 10$\times$ reduction, but after 50K gradient steps instead of 100K, since the ALOHA datasets are smaller). For the two clothes folding tasks which do not require language grounding, we still include FiLM to verify that these additional parameters do not impair task execution.
@@ A-G Additional Experiments
Ablating the OpenVLA pretrained representation. We evaluate the performance of OpenVLA-OFT policies produced by fine-tuning the underlying Prismatic VLM [20] directly on the LIBERO downstream datasets without OpenVLA’s Open X-Embodiment [33] robot pretraining. This ablation study investigates whether OpenVLA’s robot-pretrained representation remains valuable when subjected to a substantially different fine-tuning approach such as OFT. The results in Table XIV demonstrate that the variant without the pretrained OpenVLA representation consistently underperforms compared to the full OpenVLA-OFT model, confirming the benefits of using the pretrained representation for downstream policy learning.
@@
TABLE XIV: Ablation experiment: fine-tuning OpenVLA from scratch with the OFT recipe. Policy inputs here include a third-person camera image, a wrist camera image, robot proprioceptive state, and a language instruction. The from-scratch policies generally perform worse than the full OpenVLA-OFT policies, confirming that OpenVLA’s pretrained representation is beneficial for downstream policy performance even when the fine-tuning recipe differs substantially from the pretraining recipe.

### v2

@@ II Related Work
Prior works have leveraged language and vision foundation models to enhance robotic capabilities, using them as pretrained visual representations that accelerate robotic policy learning [30, 33, 31, 20, 32], for object localization in robotics tasks [9, 47], and for high-level planning and reasoning [1, 17, 44, 16, 45, 18, 6]. More recently, researchers have explored fine-tuning vision-language models (VLMs) to directly predict low-level robotic control actions, producing “vision-language-action” models (VLAs) [4, 34, 24, 23, 7, 15, 8, 53, 58, 54, 3, 2], which have demonstrated effective language following [59] and generalization to out-of-distribution test conditions and unseen semantic concepts. These works focus primarily on model development, while we focus on developing a recipe for fine-tuning such models, justifying individual design decisions with insights that we gain from our empirical analysis.
@@
Lastly, unlike prior work that relies on a separate faster low-level control policy in addition to a slower VLA [36], in this work we fine-tune the base VLA end-to-end with techniques that enable high efficiency without needing a separate controller. Further, unlike other works that collect online interaction data to continually update specialist policies via reinforcement learning [19], here we focus on the simpler imitation learning paradigm in which we develop high-performing policies just by training once offline on a fixed dataset of expert task demonstrations.
@@ IV-B Implementing Alternative Design Components
A crucial implementation detail is the choice of what represents a “feature” for modulation in vision transformers. While one might naturally consider treating individual patch embeddings as features to be modulated, we find that this approach results in poor language following. Instead, drawing from how FiLM operates in convolutional networks, where modulation applies spatially-agnostically by scaling and shifting entire feature maps, we apply each element of $\mathbf{\gamma}$ and $\mathbf{\beta}$ to the corresponding hidden unit across all visual patch embeddings so that $\mathbf{\gamma}$ and $\mathbf{\beta}$ influence all patch embeddings. Concretely, this makes $\mathbf{\gamma}$ and $\mathbf{\beta}$ $D_{ViT}$-dimensional vectors, where $D_{ViT}$ is the number of hidden dimensions (i.e., the number of elements in each of the patch embeddings in the vision transformer’s latent representations).
@@ V-A LIBERO Experimental Setup
TABLE I: LIBERO task performance results. Success rates (SR) across LIBERO benchmark task suites [26]. Results include policies fine-tuned from pretrained base models (Octo, DiT Policy, Seer, $\pi_{0}$), models trained from scratch (Diffusion Policy, Seer (scratch), MDT), and OpenVLA variants using different fine-tuning design decisions: parallel decoding (PD), action chunking (AC, chunk size $K=8$ timesteps), and continuous actions with L1 regression (Cont-L1) or with diffusion (Cont-Diffusion). Unless otherwise specified, OpenVLA diffusion policies use 50 diffusion steps at both train and test time. OpenVLA results from this work are averaged over 500 trials for each task suite (10 tasks × 50 episodes). Our full OpenVLA-OFT method with additional inputs achieves state-of-the-art 97.1% average success rate. Baseline results are from the original papers, except for Diffusion Policy, Octo, and the original OpenVLA policies, whose results are reported by Kim et al. [23]. Bold and underlined values indicate best and second-best performance. We separate the comparisons based on which inputs the policies receive as well as whether or not they were trained using modified datasets following Kim et al. [23] (in the modified versions, actions with near-zero magnitude and unsuccessful demonstrations are filtered out).   *: MDT results with 100% language annotations were obtained through direct correspondence with the authors.
@@
(Modified training dataset; unsuccessful demonstrations filtered out)
@@
OpenVLA (fine-tuned) [23] 84.7 88.4 79.2 53.7 76.5
OpenVLA (fine-tuned) + PD&AC 91.3 92.7 90.5 86.5 90.2
OpenVLA-OFT (OpenVLA (fine-tuned) + PD&AC, Cont-L1) (ours) 96.2 98.3 96.2 90.7 95.3
(Original unfiltered training dataset)
Spatial Object Goal Long Average
SR (%) SR (%) SR (%) SR (%) SR (%)
MDT (scratch; with 2% language annotations) [40] 78.5 87.5 73.5 64.8 76.1
MDT (scratch; with 100% language annotations)* 95.2 97.8 83.0
@@
OpenVLA-OFT (OpenVLA (fine-tuned) + PD&AC, Cont-L1) (ours) 95.2 94.2 95.2 93.2 94.5
@@
(Modified training dataset; unsuccessful demonstrations filtered out)
@@ V-B LIBERO Task Performance Comparisons
TABLE II: LIBERO inference efficiency results. Action generation throughput and latency for 7-dimensional actions, averaged over 100 queries on an NVIDIA A100 GPU. Each query processes a 224 × 224 px image and a LIBERO task command (“pick up the alphabet soup and place it in the basket”). We compare OpenVLA variants with parallel decoding (PD), action chunking ($K=8$ timesteps), and continuous actions using L1 regression or diffusion objectives. For the diffusion variants, we report inference efficiency measurements with varying numbers of denoising steps at test time using the DDIM sampler [46], along with the corresponding success rates in the LIBERO-Long task suite. In the final row, we report the inference efficiency of our proposed OpenVLA-OFT formulation with additional inputs, such as wrist camera image and robot state.
Throughput (Hz) $\uparrow$ Latency (Sec) $\downarrow$ LIBERO-Long SR (%)
OpenVLA 4.2 0.2396 53.7
+ PD 15.9 0.0629
+ PD&AC 108.8 0.0735 86.5
+ PD&AC, Cont-L1 109.7 0.0729 90.7
+ PD&AC, Cont-Diffusion
           $T_{train}=50$, $T_{test}=50$ 4.2 1.9070 91.1
           $T_{train}=50$, $T_{test}=10$ 19.3 0.4145 91.0
           $T_{train}=50$, $T_{test}=5$ 35.1 0.2279 90.0
           $T_{train}=50$, $T_{test}=2$ 80.3 0.0996 85.7
           $T_{train}=50$, $T_{test}=1$ 109.4 0.0731 0.0
@@
       + Additional Inputs (wrist img, proprio) 71.4 0.1120 94.5
@@ V-C LIBERO Inference Efficiency Comparisons
Results in Table II show that parallel decoding reduces latency and increases throughput by 4$\times$ by replacing 7 sequential forward passes through the decoder portion of the policy with a single pass. Adding action chunking ($K=8$) increases latency by 17% due to longer attention sequences in the decoder, but when combined with parallel decoding, it dramatically improves throughput, achieving a 26$\times$ speedup over baseline OpenVLA. The continuous actions variant with L1 regression shows negligible difference in efficiency since the additional MLP action head adds minimal computational cost compared to the base model. The primary diffusion variant requires 50 denoising steps (since this is the number of steps specified at train time) and thus suffers from high latency. However, it still achieves the same effective throughput as baseline OpenVLA due to parallel decoding and action chunking. This means that despite longer pauses between action chunks, the 50-step diffusion variant still completes robot episodes at the same speed as the original autoregressive variant. We also test the OpenVLA diffusion variant using less than 50 denoising steps at inference time (enabled by the DDIM sampler [46]), matching state-of-the-art diffusion and flow matching methods that use 5-10 steps at inference time [40, 27, 3]. These configurations result in much greater inference efficiency, as shown in Table II. However, the success rates decrease with fewer denoising steps due to reduced model quality.
@@ V-D Model Input-Output Flexibility
Evaluating these policies with additional inputs on the LIBERO benchmark reveals further improvements in average success rate across all task suites (Table I). Notably, our enhanced fine-tuned OpenVLA policies outperform even the best fine-tuned $\pi_{0}$ policies [3, 38]—which benefit from a base model with larger-scale pretraining and a more sophisticated learning objective (flow matching [25])—as well as Multimodal Diffusion Transformer (MDT) [40] and Seer [50] policies. Even with a simpler base model pretrained on less data than more recent VLAs, we find that our alternative VLA adaptation design decisions empower fine-tuned OpenVLA policies to establish a new state of the art on the LIBERO benchmark.
@@ V-F Additional Experiments
Given that the alternative fine-tuning formulation, along with additional model inputs and outputs, induces a distribution shift between the base VLA’s pretraining and fine-tuning, one reasonable question is whether the base VLA’s pretrained representations are helpful and have any influence on the results we have reported. We conduct an ablation study in Appendix A-G2 to address this question, ablating the OpenVLA pretraining phase and directly fine-tuning the underlying pretrained VLM with the OFT recipe. As shown in Table XV, the base OpenVLA pretrained representations are indeed still beneficial for robotic policy learning, as removing them leads to a 5.2% drop in average success rate (absolute) in our LIBERO evaluation suite.
(See Appendix A-G for other additional experiments, including scaling OpenVLA-OFT up to larger datasets—in both the LIBERO simulation setting as well as a real-world single-arm robot manipulation setting with over 50K demonstrations from the BridgeData V2 dataset [52].)
Figure 4: ALOHA task performance results. Comparison between policies trained from scratch (ACT, Diffusion Policy) and fine-tuned VLAs (RDT-1B, $\pi_{0}$, OpenVLA-OFT+) across four ALOHA manipulation tasks. OpenVLA-OFT+ enhances the base model with parallel decoding, action chunking, continuous actions, L1 regression, and FiLM [37] for language grounding. Fine-tuned VLAs consistently outperform from-scratch methods, with OpenVLA-OFT+ achieving the highest average performance. Scores represent average percent completion of each task (see rubrics and detailed results in Appendix A-F2).
@@ Acknowledgments
We thank the Stanford Center for Research on Foundation Models (CRFM) and Stanford Institute for Human-Centered AI (HAI) for providing computational resources that supported this research. Toyota Research Institute (TRI) provided funds to assist the authors with their research, but this article solely reflects the opinions and conclusions of its authors and not TRI or any other Toyota entity. This work was also in part supported by the Robotics and AI Institute and ONR grant N00014-22-1-2621. We thank Physical Intelligence for providing beta access to the $\pi_{0}$ model used in our ALOHA robot evaluations. We also thank Dan Fu, Karl Pertsch, and Alec Lessing for insightful discussions which contributed to the development of this work. Lastly, we are grateful to Moritz Reuss for providing additional experimental results for MDT and engaging in helpful discussions related to this work.
@@ A-D OpenVLA-OFT Hyperparameters and Training Details
OpenVLA-OFT+ training details for ALOHA. Hyperparameters for OpenVLA-OFT+ training on ALOHA tasks (with FiLM in the augmented OFT+ recipe) are shown in Table V. We maintain the same convergence criterion as in the LIBERO experiments (training until mean normalized L1 loss falls below 0.01) and similar learning rate decay strategy (again 10$\times$ reduction, but after 50K gradient steps instead of 100K, since the ALOHA datasets are smaller). For the two clothes folding tasks which do not require language grounding, we still include FiLM to verify that these additional parameters do not impair task execution.
@@ A-G Additional Experiments
A-G1 Single OpenVLA-OFT Policy for All LIBERO Task Suites Combined
In Section V and Table I, we report results with OpenVLA-OFT policies trained on each task suite independently. In this section, we assess whether our method scales to larger fine-tuning datasets by training one OpenVLA-OFT policy on all four task suites combined. As shown in Table XIV, this new policy achieves comparable average task performance as the task suite-specific policies—confirming that our method scales to larger fine-tuning datasets.
A-G2 Ablating FiLM in LIBERO
The FiLM ablation study in Section VI suggests that FiLM is crucial for enabling strong language following in real-world ALOHA robot tasks. In this section, we assess whether FiLM is similarly important in the LIBERO simulation tasks. We train a single OpenVLA-OFT+ policy (with FiLM) on all LIBERO task suites combined (similar to the previous section) and report task performance results in Table XIV, comparing performance against the OpenVLA-OFT policy trained without FiLM. In LIBERO, FiLM leads to slightly higher average success rate, though the difference here is minor compared to the findings in the real-world ALOHA experiments, where language following is more challenging due to the reasons discussed in Section IV-C.
A-G3 Ablating the OpenVLA Pretrained Representation
We evaluate the performance of OpenVLA-OFT policies produced by fine-tuning the underlying Prismatic VLM [21] directly on the LIBERO downstream datasets without OpenVLA’s Open X-Embodiment [34] robot pretraining. This ablation study investigates whether OpenVLA’s robot-pretrained representation remains valuable when subjected to a substantially different fine-tuning approach such as OFT. The results in Table XV demonstrate that the variant without the pretrained OpenVLA representation consistently underperforms compared to the full OpenVLA-OFT model, confirming the benefits of using the pretrained representation for downstream policy learning.
A-G4 Scaling Up OpenVLA-OFT to a Larger Real-World Dataset (BridgeData V2)
In Appendix A-G1, we observe that a single OpenVLA-OFT policy can effectively fit all four LIBERO task suite datasets combined, confirming that the proposed method scales to larger fine-tuning datasets. In this section, we scale up the fine-tuning data further and assess whether OpenVLA-OFT can also fit a real-world robotic manipulation dataset that is significantly larger and more diverse than both the LIBERO datasets and the ALOHA robot datasets discussed in Section VI. Specifically, we train OpenVLA-OFT (without FiLM) on the BridgeData V2 dataset [52], which contains 50,365 real WidowX robot demonstrations (25$\times$ more than the four LIBERO task suites combined).Although the base OpenVLA model was already pretrained on Bridge data, we note that the model must still be fine-tuned when using the new OFT recipe since the architecture, learning algorithm, and action representation and decoding scheme have been modified significantly from the pretraining setup. In fact, the initial action regression L1 loss in the beginning of OpenVLA-OFT training on Bridge is large (roughly 0.5 when actions are normalized to the scale $[-1,+1]$).
We evaluate OpenVLA-OFT on a subset of BridgeData V2 WidowX robot tasks from the evaluation suite used in the original OpenVLA work [23]. This representative subset covers the four types of generalization (visual, motion, physical, semantic) as well as language grounding tasks. We compare the task performance of OpenVLA-OFT with that of the public OpenVLA checkpoint, scoring both methods using the same criteria used in the OpenVLA work. As shown in Table XVI, OpenVLA-OFT surpasses OpenVLA on average across these tasks (69.2% versus 65.8%, respectively)—confirming scalability of the proposed OFT approach to much larger and more diverse real robot datasets. Further, we observe that FiLM is not necessary for satisfactory language following in Bridge tasks, likely because the base model has shown effective language following in Bridge tasks.
@@ A-G4 Scaling Up OpenVLA-OFT to a Larger Real-World Dataset (BridgeData V2)
TABLE XIV: Additional LIBERO experiments: Single policy for all LIBERO task suites and FiLM ablation study. Here we train a single policy on all four LIBERO task suites combined, without FiLM (“OpenVLA-OFT”) and with FiLM (“OpenVLA-OFT+”). Policy inputs here include a third-person camera image, wrist camera image, robot proprioceptive state, and a language instruction. We observe comparable task performance when training on all task suites together versus training on each of them independently. In addition, we find that FiLM has little effect on task performance in LIBERO, given that the variant without FiLM already exhibits good language following.
Spatial Object Goal Long Average
SR (%) SR (%) SR (%) SR (%) SR (%)
OpenVLA-OFT (1 policy per suite) (original, from Table I) 97.6 98.4 97.9 94.5 97.1
OpenVLA-OFT (1 policy for all 4 suites) 97.7 98.0 96.1 95.3 96.8
OpenVLA-OFT+ (1 policy for all 4 suites, + FiLM) 97.8 98.2 98.2 93.8 97.0
TABLE XV: Ablation experiment: Fine-tuning OpenVLA from scratch with the OFT recipe. Policy inputs here include a third-person camera image, a wrist camera image, robot proprioceptive state, and a language instruction. The from-scratch policies generally perform worse than the full OpenVLA-OFT policies, confirming that OpenVLA’s pretrained representation is beneficial for downstream policy performance even when the fine-tuning recipe differs substantially from the pretraining recipe.
@@
TABLE XVI: OpenVLA-OFT in BridgeData V2 WidowX robot evaluations. We use a subset of the tasks from the original OpenVLA BridgeData V2 evaluation suite, along with the corresponding scoring criteria. Each score is the average over 10 trials per task. OpenVLA-OFT surpasses OpenVLA in average task performance and even demonstrates effective language following without FiLM.
Category Task OpenVLA Score OpenVLA-OFT Score
Visual gen Put Eggplant into Pot 60 90
Motion gen Lift Eggplant 70 60
Physical gen Flip Pot Upright 90 80
Semantic gen Lift White Tape 0 20
Language grounding Put {Blue Cup, Pink Cup} on Plate 85 90
Language grounding Lift {Cheese, Red Chili Pepper} 90 75
Average 65.8 69.2